OpenStudy (anonymous):

There are 50 cans of soda in a cooler. The table shows the number of each kind of soda. Two cans of soda are taken from the cooler at random. What is the probability that both cans are Lemon? A. 3/35 B. 9/100 C. 41/7 D. 3/5

OpenStudy (anonymous):

I got A as an answer

OpenStudy (agent0smith):

Total is 50. First can, there's 15 lemon: 15/50 = 3/10 second can, 14 lemon and 49 remaining cans: 14/49 = 2/7 multiplying: 3/10 * 2/7 = 3/35

OpenStudy (yttrium):

SWAG.

OpenStudy (anonymous):

LOL!